Solve for x.
1+ 10 over x + 26 over x^2 = 0
i could not write this out so i had to spell it out, i hope you can understand this.

Answer by stanbon(75887)

1+ 10 over x + 26 over x^2 = 0
------
1 + (10/x) + (26/x^2) = 0
-------
Multiply thru by x^2 to get:
x^2 + 10x + 26 = 0
x = [-10 +- sqrt(100-4*26)]/2
----
x = [-10 +- sqrt(-4)]/2
x = [-10 +- 2i]/2
----
x = -5+i or x = -5-i
==========================
Cheers,
Stan H.